OpenStudy (anonymous):

Use the distributive property to simplify the expression below -3x -9 / -3

OpenStudy (zepp):

@malcria16 I've already answered this question.

OpenStudy (anonymous):

\[-(3x+9)\div-3=(3x+9)\div3=x+3\]